A 2.0 kg particle moves as a function of time as follows:

             x = 3.5 sin (1.5t+π/2)

Where, distance is measured in meters and time in seconds.

(a) What is the amplitude, frequency, angular frequency, and period of this motion?



1
Expert's answer
2021-05-07T09:44:56-0400

Amplitude "A=3.5 \ (m)"


Angular frequency "\omega=1.5\ (rad/s)"


Frequency "f=\frac{\omega}{2\pi}=\frac{1.5}{2\cdot 3.14}=0.24\ (Hz)"


Period "T=\frac{1}{f}=\frac{1}{0.24}=4.17\ (s)"